# Human foraging strategies flexibly adapt to resource distribution and time constraints

### v3

@@ 4 Discussion
Concerning the first question, we found that that participants showed flexibility in foraging behaviour and were sensitive to both resources distribution and time availability. While some results (e.g., more coins collected with longer batteries) are simply the expected effect of our manipulations, others are indicatives of strategy. A key indication of adaptive foraging strategies comes from the analysis of boxes opened in the first area. Participants collected more boxes in the first area when doing so was more favourable, namely, when they were in a rich environment and their battery duration was longer (Figure 3). Interestingly, when their first visited area was rich, they opened more boxes when the environment was mixed than when it was rich, showing sensitivity to these different conditions (Figure S2). Furthermore, participants modulated their strategy and their effort as a function of both richness of the environment and foraging time. First, participants opened more boxes in rich environments, suggesting that they exploited more and potentially invested more effort with more points to be earned. Second, and more remarkably, participants showed faster navigation time between boxes within the same area when foraging time was shorter. This suggests that with more limited time available, participants exerted more effort, analogous to previous findings (Lewis et al.,, 2004), but did not show less efficient foraging (Wu et al.,, 2022). Taken together, these results show that participants flexibly adapt various aspects of their foraging strategy – including the choice of when to leave areas and how fast to move – to resource distribution and time availability. Future studies might assess more fine-grained aspects of adaptive foraging strategy, by manipulating other aspects of the task, such as the uncertainty about rewards and their volatility.
@@
Concerning the second question, we found that during the experiment, participants showed a clear learning trend and an overall increase of performance, as indexed for example by the number of coins collected (Figure 5.a). The increase of performance was plausibly due to the reduction of uncertainty about resource locations and distributions – and to a combination of improved skill and cognitive strategy. First, during the experiment participants learned how to navigate more effectively and what are the most likely locations of boxes, and became faster in locating and opening them (Figure 5.f). Second, participants tended to modify their strategy, from more exploitative to more exploratory, by reducing the number of boxes opened in the first area (Figure 5.b) – plausibly, as they become more confident that they can effectively locate boxes across all areas. Our results suggest a transition between an initial, more exploratory phase in which participants need to learn task statistics to a subsequent phase in which participants develop and exploit a more effective strategy that approximates an optimal forager without uncertainty, but without fully reaching it (see below). It is however worth noting that while uncertainty can be decreased during the task, it cannot be eliminated; for example, the location of the rich areas in “mixed environments” remains unknown and requires exploration until the end of the experiment. In “mixed environments”, it is only after exploring two areas that a participant can be sure about which areas are rich – but at this point, the trial would be almost finished. Furthermore, the locations of the boxes and the time needed to locate and reach them changes from trial to trial, making fixed strategies (e.g., always collect two boxes before leaving an area) less effective. This implies that despite uncertainty decreases during the experiment, a complete passage from exploration to exploitation is impossible and participants’ foraging behavior has to remain flexible across all the experiment. We also found that the order of blocks experienced by participants showed two significant effects. First, participants who performed the long battery block in the first part of the experiment showed a better performance in the second half of the experiment, when they performed the short battery block, compared to participants who performed the short battery block first. This finding suggests that the longer training time enjoyed by participants who performed the long battery block first translates into an advantage in the second block. Second, participants who started the experiment with the short battery block visited significantly more areas in the second (long battery) block. A possible explanation is that during the short battery block, participants learned it was more convenient to leave areas quickly to access the more rewarding boxes and this behavior persisted in the second experimental block. Taken together, these two effects indicate that training time and strategies learned during the fist block influences subsequent performance, highlighting the importance of studying learning dynamics and not only steady-state behavior during foraging experiments. Future studies could assess what sequence of blocks (or training curriculum) could be more advantageous to improve foraging strategies.
@@
Concerning the third question, we found that participants’ performance – in terms of total coins collected – differed significantly from an optimal agent in the trials with long battery, but not in those with short battery. The fact that participants in short battery trials approximated better optimal solutions could be due to a combination of (at least) two factors, one more strategic (as discussed above, participants are faster in short battery trials, see Figure 6) and another purely statistical (the fact that both participants and the optimal agent collect less coins in short battery trials and hence their differences might not reach statistical significance). Participants differed significantly from the optimal agent also for the number of boxes collected in the first area (which indexes the flexible adaptation of foraging strategies), during both long battery and short battery trials. While the difference between participants and the optimal agent was not visible when looking at the results of the short battery trials in the mixed environment (Figure 7), further splitting the mixed environment data into trials in which the first visited area was rich or poor shows that in the former case participants underharvest, whereas in the latter case they overharvest (Figure 8). The differences between the participants and the optimal agent are not surprising, especially when considering that the optimal agent has perfect knowledge of the environment, whereas participants make decisions under various sources of uncertainty, such as where and when they will be able to find the next box (at at least in the first part of the experiment, their expected content). These differences might translate into more or less deterministic choices. For example, the results of the mixed environment (Figure 8) showed that the optimal agent privileged extreme solutions, to either collect all the five available boxes (if the area is rich and the battery long) or to collect just one box and leave (if the area is poor and the battery short). Rather, participants selected similar but less extreme solutions, showing some “regression to the mean”, which could be explained by participants’ uncertainty and/or additional stochasticity in their decision processes. Additionally, the tendency to overharvest in poor areas may be partially driven by the sunk cost effect (Arkes and Blumer,, 1985), where individuals persist in a behavior due to prior investments, even when it is suboptimal. Finally, we found that over the course of the experiment, participants’ performance approximated but did not reach the optimal agent (Figure S4), suggesting that even with more training, participants’ strategies retain some aspects that are not well captured by optimal agents with perfect knowledge. These differences between participants’ and optimal agents’ strategies remain to be investigated in future studies.
@@
One limitation of the current study is that we examined foraging decisions using a relatively simple model. This approach was taken because we aimed to investigate the extent to which people approximate optimal foraging strategies during the experiment, assuming no uncertainty except for the location of rich and poor areas. Future studies could formalize alternative strategies that individuals might adopt when deciding how to forage and when to leave a patch. Previous research has identified possible foraging rules (e.g., giving-up time and area-restricted search in (Wilke et al.,, 2009)) that could be adapted to our task and used for model comparison, providing further insight into how people forage and whether their strategies change between early and later phases of foraging. Additionally, incorporating learning models could offer a more detailed examination of explore-exploit dynamics during foraging (Friston et al.,, 2016; Sutton,, 2018). Finally, future studies might adapt our setup to clinical or subclinical populations to investigate under which conditions they develop adaptive or maladaptive foraging strategies (Addicott et al.,, 2017; Barack et al.,, 2024; Pellicano et al.,, 2011).
@@
In sum, our results highlight the great flexibility of human foraging in response to varying resource distributions, limited foraging time, and learning opportunities. Furthermore, they showcase the importance of studying foraging in rich environments that engage various cognitive processes, such as orientation and spatial navigation, while also affording learning and strategy adaptation over time. These ecologically valid contexts enable new questions about how foraging strategies adapt to the physical properties of the environment, revealing subtle aspects often overlooked in accounts like optimal foraging theory. For example, our experiment suggests that people transition from an initial exploration phase—where they reduce uncertainty about key aspects of the task, such as resource locations and travel times—to a subsequent exploitation phase, in which they leverage their knowledge to deploy more effective foraging strategies. These learning and adaptation effects in foraging present a valuable avenue for future research.
Acknowledgements
@@
This research received funding from the European Union’s Horizon 2020 Framework Programme for Research and Innovation under the Specific Grant Agreements No. 952215 (TAILOR) to G.P.; the European Research Council under the Grant Agreement No. 820213 (ThinkAhead) to G.P.; the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union – NextGenerationEU (Project IR0000011, CUP B51E22000150006, “EBRAINS-Italy”; Project PE0000013, CUP B53C22003630006, “FAIR”; Project PE0000006, CUP J33C22002970002 “MNESYS”) to G.P., PRIN PNRR P20224FESY to G.P., and the National Research Council, project iForagers. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript. We used a Generative AI model to correct typographical errors and edit language for clarity.
@@ A.4 Supplementary Figures
Figure S3: Participants’ average time to leave a room after collecting the last treasure box within it, across different experimental conditions. The red line represents the linear regression fit, highlighting a significant decrease in time to exit as participants progressed through the trials (R² = 0.19, p = 0.005). See the main text for explanation.

### v4

@@ 4 Discussion
Concerning the first question, we found that that participants showed flexibility in foraging behaviour and were sensitive to both resources distribution and time availability. While some results (e.g., more coins collected with longer batteries) are simply the expected effect of our manipulations, others are indicatives of strategy. A key indication of adaptive foraging strategies comes from the analysis of boxes opened in the first area. Participants collected more boxes in the first area when doing so was more favourable, namely, when they were in a rich environment and their battery duration was longer (Figure 3). Interestingly, when their first visited area was rich, they opened more boxes when the environment was mixed than when it was rich, showing sensitivity to these different conditions (Figure S2). Furthermore, participants modulated their strategy and their effort as a function of both richness of the environment and foraging time. First, participants opened more boxes in rich environments, suggesting that they exploited more and potentially invested more effort with more points to be earned. Second, and more remarkably, participants showed faster navigation times between boxes within the same area when foraging time was shorter. Interestingly, this occurred not only in rich (and less uncertain) environments but also in mixed (and more uncertain) ones. In short battery trials within mixed environments, participants necessarily gathered less information than in long battery trials, leading to greater uncertainty—something that would typically call for more, not less, deliberation and slower movement. The fact that participants were faster rather than slower in short battery trials suggests that limited time prompted greater effort, consistent with previous findings (Lewis et al.,, 2004), without resulting in less efficient foraging (Wu et al.,, 2022). This aligns with recent work by (Pisauro et al.,, 2024), which showed that under deadline pressure, effort can be perceived as a valued means to achieve goals rather than merely a cost to minimize. Taken together, these results show that participants flexibly adapt various aspects of their foraging strategy – including the choice of when to leave areas and how fast to move – to resource distribution and time availability. Future studies might assess more fine-grained aspects of adaptive foraging strategy, by manipulating other aspects of the task, such as the uncertainty about rewards and their volatility.
@@
Concerning the second question, we found that during the experiment, participants showed a clear learning trend and an overall increase of performance, as indexed for example by the number of coins collected (Figure 5.a). The increase of performance was plausibly due to the reduction of uncertainty about resource locations and distributions – and to a combination of improved skill and cognitive strategy. First, during the experiment participants learned how to navigate more effectively and what are the most likely locations of boxes, and became faster in locating and opening them (Figure 5.f). Second, participants modified their patch leaving strategy across trials. Specifically, they reduced their initial overharvesting and the number of boxes opened in the first area (Figure 5.b), as they become more confident that they can effectively locate boxes across all the other areas. Our results suggest a transition between an initial, more exploratory phase in which participants need to learn task statistics to a subsequent phase in which participants develop and exploit a more effective strategy that approximates an optimal forager without uncertainty, but without fully reaching it (see below). It is, however, worth noting that while uncertainty decreases during the task, it cannot be fully eliminated. For example, the locations of the boxes and the time needed to locate and reach them vary from trial to trial, making fixed strategies (e.g., always collecting two boxes before leaving an area) less effective. Furthermore, in “mixed environments,” it is only after exploring two areas that a participant can be certain about which areas are rich. Consistent with this, we found that within-trial uncertainty influences participants’ behavior and deliberation time. In particular, the observed reduction in exit time across visited areas in mixed environments suggests that participants deliberated more when uncertainty was higher and accelerated their decisions once they had gathered enough information to reduce ambiguity about area richness. Notably, this pattern was absent in rich environments, where full information about resource richness was provided at the outset—supporting our interpretation that the effect is driven by uncertainty reduction rather than motor or attentional factors. We also found that the order of blocks experienced by participants showed two significant effects. First, participants who performed the long battery block in the first part of the experiment showed a better performance in the second half of the experiment, when they performed the short battery block, compared to participants who performed the short battery block first. This finding suggests that the longer training time enjoyed by participants who performed the long battery block first translates into an advantage in the second block. Second, participants who started the experiment with the short battery block visited significantly more areas in the second (long battery) block. A possible explanation is that during the short battery block, participants learned it was more convenient to leave areas quickly to access the more rewarding boxes and this behavior persisted in the second experimental block. Taken together, these two effects indicate that training time and strategies learned during the fist block influences subsequent performance, highlighting the importance of studying learning dynamics and not only steady-state behavior during foraging experiments. Future studies could assess what sequence of blocks (or training curriculum) could be more advantageous to improve foraging strategies and whether, with sufficient time, human participants could match the performance of optimal agents.
@@
Concerning the third question, we found that participants’ performance – in terms of total coins collected – differed significantly from an optimal agent in the trials with long battery, but not in those with short battery. The fact that participants in short battery trials approximated better optimal solutions could be due to a combination of (at least) two factors: first, as discussed above, participants are faster in short battery trials; second, both participants and the optimal agent collect fewer coins in short battery trials and the difference might not rise statistical significance. Participants differed significantly from the optimal agent also for the number of boxes collected in the first area (which indexes the flexible adaptation of foraging strategies), during both long battery and short battery trials. While the difference between participants and the optimal agent was not visible when looking at the results of the short battery trials in the mixed environment (Figure 7), further splitting the mixed environment data into trials in which the first visited area was rich or poor shows that in the former case participants underharvest, whereas in the latter case they overharvest (Figure 8). The differences between the participants and the optimal agent are not surprising, especially when considering that the optimal agent has perfect knowledge of the environment, whereas participants make decisions under various sources of uncertainty, such as where and when they will be able to find the next box (at at least in the first part of the experiment, their expected content). These differences might translate into more or less deterministic choices. For example, the results of the mixed environment (Figure 8) showed that the optimal agent privileged extreme solutions, to either collect all the five available boxes (if the area is rich and the battery long) or to collect just one box and leave (if the area is poor and the battery short). Rather, participants selected similar but less extreme solutions, showing some “regression to the mean”, which could be explained by participants’ uncertainty and/or additional stochasticity in their decision processes. Finally, we found that over the course of the experiment, participants’ performance approximated but did not reach the optimal agent (Figure S5), suggesting that even with more training, participants’ strategies retain some aspects that are not well captured by optimal agents with perfect knowledge. Understanding what causes the differences between participants’ and optimal agents’ strategies (e.g., heuristics against “extreme” solutions or some rational trade-off) remains to be investigated in future studies.
@@
One limitation of the current study is that we examined foraging decisions using a relatively simple model, which assumed uncertainty only about the locations of rich and poor areas. This approach was chosen because we aimed to investigate how closely people approximate optimal foraging strategies during the experiment, as they progressively learned the task structure—and thus, ideally, remained uncertain only about the locations of rich and poor areas, similar to the optimal model. Future studies could formalize alternative strategies that individuals might adopt when deciding how to forage and when to leave a patch. Previous research has identified possible foraging rules (e.g., giving-up time and area-restricted search in (Wilke et al.,, 2009)) that could be adapted to our task and used for model comparison, providing further insight into how people forage and whether their strategies change between early and later phases of foraging. Additionally, incorporating learning models could offer a more detailed examination of explore-exploit dynamics during foraging (Friston et al.,, 2016; Sutton,, 2018). Moreover, despite the present study focuses on behavior, our paradigm was informed by neuroscientific models of foraging that implicate brain areas such as the dorsal anterior cingulate cortex and the hippocampus in regulating stay-or-leave decisions, and dopaminergic systems in effort allocation under time constraints (Wittmann et al.,, 2016; Le Heron et al.,, 2019; Ianni et al.,, 2023; Fine et al.,, 2025). Future work could adapt our task to investigate the neural processes supporting flexible foraging with varying resource distribution and limited time. Finally, future studies might adapt our setup to clinical or subclinical populations to investigate under which conditions they develop adaptive or maladaptive foraging strategies (Addicott et al.,, 2017; Barack et al.,, 2024; Pellicano et al.,, 2011).
@@
In sum, our results highlight the great flexibility of human foraging in response to varying resource distributions, limited foraging time, and learning opportunities. Furthermore, they showcase the importance of studying foraging in rich environments that engage various cognitive processes, such as orientation and spatial navigation, while also affording learning and strategy adaptation over time. These ecologically valid contexts enable new questions about how foraging strategies adapt to the physical properties of the environment, revealing subtle aspects often overlooked in accounts like optimal foraging theory. For example, our experiment suggests that people transition from an initial exploration (or information gathering) phase—where they reduce uncertainty about key aspects of the task, such as resource locations and travel times—to a subsequent exploitation (or reward harvesting) phase, in which they leverage their knowledge to deploy more effective foraging strategies. These learning and adaptation effects in foraging present a valuable avenue for future research.
Declarations
Funding
@@
This research received funding from the European Union’s Horizon 2020 Framework Programme for Research and Innovation under the Specific Grant Agreements No. 952215 (TAILOR) to G.P.; the European Research Council under the Grant Agreement No. 820213 (ThinkAhead) to G.P.; the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union – NextGenerationEU (Project IR0000011, CUP B51E22000150006, “EBRAINS-Italy”; Project PE0000013, CUP B53C22003630006, “FAIR”; Project PE0000006, CUP J33C22002970002 “MNESYS”) to G.P., PRIN PNRR P20224FESY to G.P., and the National Research Council, project iForagers. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
Conflicts of interest/Competing interests
The authors have declared that no competing interests exist.
Ethics approval
All the procedures were approved by the CNR Ethics committee.
Consent to participate
Informed consent was obtained from all participants prior to their inclusion in the study. Participants were provided with detailed information about the study’s purpose, procedures, and potential risks, and they voluntarily agreed to participate.
Consent for publication
Informed consent for publication was obtained from all participants. They were provided with detailed information about the study and its intended dissemination, and they voluntarily agreed to the publication of the findings.
Availability of data and materials
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Code Availability
Scripts used for the main model, statistical analysis, and result visualization are provided at the following link: https://github.com/DavideNuzzi/Human-foraging-strategies-flexibly-adapt-to-resource-distribution-and-time-constraints
Authors’ contributions
All authors contributed to the research and manuscript preparation. Valeria Simonelli was responsible for Conceptualization, Methodology, Data Collection, Data Curation, Formal Analysis, Validation, Visualization, and Writing – Original Draft and Review. Davide Nuzzi contributed to Methodology, Data Curation, Formal Analysis, Validation, Software, Visualization, and Writing – Original Draft and Review. Gian Luca Lancia was involved in Conceptualization, Methodology, and Writing – Review. Giovanni Pezzulo contributed to Conceptualization, Methodology, Supervision, Writing – Original Draft and Review, and Funding Acquisition.
@@ Appendix A Supplementary analyses
Environment Battery Contrast Estimate Std Error p-value
Mixed . Short - Long -28.356 0.673 $<0.001$
Rich . Short - Long -33.956 0.673 $<0.001$
. Short Mixed - Rich -7.241 0.673 $<0.001$
. Long Mixed - Rich -12.841 0.673 $<0.001$
Table S7: Pairwise post-hoc comparisons for the variable: total coins
Environment Battery Contrast Estimate Std Error p-value
Mixed . Short - Long -0.600 0.081 $<0.001$
Rich . Short - Long -0.794 0.081 $<0.001$
. Short Mixed - Rich -0.318 0.081 $<0.001$
. Long Mixed - Rich -0.512 0.081 $<0.001$
Table S8: Pairwise post-hoc comparisons for the variable: boxes collected in first area
Environment Battery Contrast Estimate Std Error p-value
Mixed . Short - Long -5.432 0.113 $<0.001$
Rich . Short - Long -5.415 0.113 $<0.001$
. Short Mixed - Rich -0.244 0.113 0.118
. Long Mixed - Rich -0.226 0.113 0.170
Table S9: Pairwise post-hoc comparisons for the variable: boxes collected
Environment Battery Contrast Estimate Std Error p-value
Mixed . Short - Long 0.454 0.032 $<0.001$
Rich . Short - Long 0.452 0.032 $<0.001$
. Short Mixed - Rich -0.999 0.032 $<0.001$
. Long Mixed - Rich -1.001 0.032 $<0.001$
Table S10: Pairwise post-hoc comparisons for the variable: average coins per box
Environment Battery Contrast Estimate Std Error p-value
Mixed . Short - Long -0.765 0.044 $<0.001$
Rich . Short - Long -0.750 0.044 $<0.001$
. Short Mixed - Rich 0.050 0.044 0.695
. Long Mixed - Rich 0.065 0.044 0.459
Table S11: Pairwise post-hoc comparisons for the variable: number of visited areas
Environment Battery Contrast Estimate Std Error p-value
Mixed . Short - Long -0.263 0.073 0.001
Rich . Short - Long -0.249 0.073 0.003
. Short Mixed - Rich 0.022 0.073 0.997
. Long Mixed - Rich 0.036 0.073 0.980
Table S12: Pairwise post-hoc comparisons for the variable: time between boxes
@@
First Area Battery Contrast Estimate Std Error p-value
Poor . Short - Long -29.008 0.836 $<0.001$
Rich . Short - Long -27.663 0.797 $<0.001$
. Short Poor - Rich -1.931 0.822 0.074
. Long Poor - Rich -0.586 0.825 0.926
Table S19: Pairwise post-hoc comparisons for the variable: total coins
First Area Battery Contrast Estimate Std Error p-value
Poor . Short - Long -0.290 0.100 0.016
Rich . Short - Long -0.769 0.095 $<0.001$
. Short Poor - Rich -1.205 0.098 $<0.001$
. Long Poor - Rich -1.684 0.099 $<0.001$
Table S20: Pairwise post-hoc comparisons for the variable: boxes collected in first area
First Area Battery Contrast Estimate Std Error p-value
Poor . Short - Long -5.404 0.159 $<0.001$
Rich . Short - Long -5.460 0.152 $<0.001$
. Short Poor - Rich 0.047 0.156 0.997
. Long Poor - Rich -0.009 0.157 1.000
Table S21: Pairwise post-hoc comparisons for the variable: boxes collected
First Area Battery Contrast Estimate Std Error p-value
Poor . Short - Long 0.263 0.053 $<0.001$
Rich . Short - Long 0.648 0.051 $<0.001$
. Short Poor - Rich -0.455 0.052 $<0.001$
. Long Poor - Rich -0.069 0.052 0.560
Table S22: Pairwise post-hoc comparisons for the variable: average coins per box
First Area Battery Contrast Estimate Std Error p-value
Poor . Short - Long -0.718 0.064 $<0.001$
Rich . Short - Long -0.829 0.061 $<0.001$
. Short Poor - Rich 0.326 0.063 $<0.001$
. Long Poor - Rich 0.215 0.063 0.003
Table S23: Pairwise post-hoc comparisons for the variable: number of visited areas
First Area Battery Contrast Estimate Std Error p-value
Poor . Short - Long -0.291 0.109 0.030
Rich . Short - Long -0.223 0.104 0.121
. Short Poor - Rich -0.202 0.107 0.217
. Long Poor - Rich -0.134 0.107 0.615
Table S24: Pairwise post-hoc comparisons for the variable: time between boxes
@@ A.4 Supplementary Figures
Figure S3: Number of coins collected by participants before leaving (a) rich or (b) poor areas. The number of collected coins changes significantly across conditions ($p=0.041$ for the environment richness comparison in the short battery condition, $p<0.001$ in all the other cases, Mann-Whitney U-test, FDR correction), suggesting that participants do not adopt a threshold-driven strategy to decide when to leave areas. See the main text for explanation.
Figure S4: Participants’ exit time: average time to leave a room after collecting the last treasure box within it, across different experimental conditions. The red line represents the linear regression fit, highlighting a significant decrease in time to exit as participants progressed through the trials (R² = 0.19, p = 0.005). See the main text for explanation.